Help In the country of Economica, the total labor force consists of 10,000 workers 600 of these workers are unemployed, and thus the current unemployment rate is 6 0% This year, 200 workers will lose their jobs, 170 workers will return to the labor force after an absence. 100 workers will voluntarily leave their jobs, and 500 new workers will enter the labor force At the same time, 830 workers will find jobs, and 300 workers will retire or leave the labor force for other reasons Thus, at the end of the year, the labor force will consist ofworkers, and the unemployment rate will be Round your espo sest one de mal pace
Answer #1

Assumption - we are assuming that the workers who leave their job voluntarily will still be looking for another job and thus will not be excluded from labor force, we could have assumed otherwise in that case instead of adding these 100 worker in umemployed workers category we would have deducted them from Labor force

Labor force at the end of year = 10000+170 (return to labor force)+ 500 (new workers) - 300 (retire or leave labor force) = 10370

Umemployed workers = 600+200 (lose jobs)+170 (return to labor force)+100 (leaving the jobs voluntarily)+ 500 (new workers) - 830 (new job found) = 740

Thus at the end of the year labor force will consist of 10370 workers and unemployment rate will be (740/10370) 7.135969% i.e. 7.1%